Circular No. 8823

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 F1 (LONEOS) An apparently asteroidal object discovered by the LONEOS project

(discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found to be cometary by several CCD astrometrists. J. Young (Table Mountain, 0.61-m f/16 Cassegrain reflector) found the object to have a slightly elongated (north-to-south), 8′′ coma with no central condensation and a hint of a short, straight 12′′ tail in p.a. 175^{o} on images taken on Mar. 19.5 UT; his images from Mar. 20.22–20.25, taken through very light cirrus clouds, confirm a 12′′ tail in p.a. 185^{o} and a 4′′ coma with some condensation. Twenty-six co-added 60-s unfiltered images taken by L. Donato, M. Gonano, V. Santini, and G. Sostero (Remanzacco, Italy, 0.45-m f/4.4 reflector) on Mar. 20.9 reveal the presence of a compact coma of diameter almost 20′′ and total mag near 18.5. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 Schmidt-Cassegrain reflector) writes that his co-added exposures (totalling 22.5 min) from Mar. 21.1 show a thin, straight tail 90′′ long in p.a. 235^{o}, and a coma diameter of 7′′.

2007 UT  α_{2000}  δ_{2000}  Mag.
Mar. 19.26399  11^{h}22^{m}57.^{s}60  +20^{o}11′58′′.2  19.5

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-F52.

T = 2007 Nov. 3.288 TT  ω = 152.^{o}188
Ω = 173.522 }2000.0
q = 0.44459 AU  i = 118.707

SUPERNOVAE 1996cr AND 2004ip An apparent type-IIn supernova initially identified as an x-ray source

in ESO 97-G13, and subsequently identified by F. Bauer and colleagues on numerous archival CCD images obtained with the Anglo-Australian Telescope in mid-March 1996, has been given the designation SN 1996cr (details on CBET 879). An apparent supernova reported by S. Mattila et al. from K_{s}-band infrared images in Sept. 2004, and designated then as PSN K0409-001 (cf. IAUC 8335), has been now designated as SN 2004ip, with additional details given on CBET 858.

2007 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green